OpenStudy (anonymous):

for the pair of functions, f(x)=x+2 and g9x)=x^2+x-3, find the following a).(f*g)(x) b).(f*g)(4) c).(g*f)(x) d).(g*f)(4)

OpenStudy (anonymous):

(f*g)(x)=f(x)g(x)=(x+2)(x^2+x-3) I'm sure you can do the rest on your own.

OpenStudy (anonymous):

a. (x^2+x-3)+2 = x^2+x-1 b. (4^2+4-3)+2 = 19 c. (x+2)^2+(x+2)-3 = x^2+4x+4+x+2-3 = x^2 +5x+3 d. (4)^2+5(4)+3 =39

OpenStudy (anonymous):

That's not true, right? He asked for (f*g)(x), not f(g(x)).